Mika has three and seven-sixteenths cm of wire.she needs only two and five-eigths cm of wire for her prodject in TLE how much wire does she need to cut??


1
Expert's answer
2020-10-26T19:57:24-0400

Length of wire Mika has"= 3\frac{7}{16}cm = \frac{55}{16}cm"


Length of wire Mika needs "= 2\frac{5}{8}cm = \frac{21}{8}cm"


Length she has = Length she needs + Length she cuts


"\therefore" Length she cuts = Length she has "-" Length she needs


Length she cuts "=\frac{55}{16} - \frac{21}{8} = \frac{13}{16}cm"


"\therefore" The length of wire she cuts is thirteen-sixteenths cm